OpenStudy (anonymous):

Sam decides to build a square garden. If the area of the garden is 25x2 + 30x + 9 square feet, what is the length of one side of the garden? (3x + 5) feet (3x − 5) feet (5x + 3) feet (5x − 3) feet

Parth (parthkohli):

Hint: One side is square root of the area.

Parth (parthkohli):

Find\[\sqrt{25x^2 + 30x + 9} \]

OpenStudy (anonymous):

square the given options.. whichever gives you the area.. is the right choice..

Parth (parthkohli):

\[25x^2 +30x + 9 \implies 25x^2 + 15x + 15x + 9 \implies 5x(5x + 3) + 3(5x + 3) \]\[ \implies (5x + 3)(5x + 3) \implies (5x + 3)^2\]

Parth (parthkohli):

What is \(\sqrt{(5x+ 3)^2}\)

OpenStudy (anonymous):

I CANT FIND MY CALCULATOR!!!

Parth (parthkohli):

No need to.\[\sqrt{number ^2}=number \]

OpenStudy (anonymous):

-_- oh... hehe

OpenStudy (anonymous):

so its the third one?